Use appropriate algebra and Theorem 7.2.1 to fin s L { s2 + 3s + 3s - 4 Х Need Help? Read It Talk to a Tutor
Answer #1

we have

L^{-1}\left \{ \frac{s}{s^2+3s-4} \right \}=L^{-1}\left \{ \frac{s}{(s-1)(s+4)} \right \}

L^{-1}\left \{ \frac{s}{s^2+3s-4} \right \}=\frac{1}{5}L^{-1}\left \{ \frac{5s}{(s-1)(s+4)} \right \}

L^{-1}\left \{ \frac{s}{s^2+3s-4} \right \}=\frac{1}{5}L^{-1}\left \{ \frac{s+4+4s-4}{(s-1)(s+4)} \right \}

L^{-1}\left \{ \frac{s}{s^2+3s-4} \right \}=\frac{1}{5}L^{-1}\left \{ \frac{(s+4)+4(s-1)}{(s-1)(s+4)} \right \}

L^{-1}\left \{ \frac{s}{s^2+3s-4} \right \}=\frac{1}{5}L^{-1}\left \{ \frac{1}{s-1}+\frac{4}{s+4} \right \}

L^{-1}\left \{ \frac{s}{s^2+3s-4} \right \}=\frac{1}{5}L^{-1}\left \{ \frac{1}{s-1} \right \}+\frac{4}{5}L^{-1}\left \{ \frac{1}{s+4} \right \}

we know that,

L^{-1}\left \{ \frac{1}{s-a} \right \}=e^{at}

we can say that,

L^{-1}\left \{ \frac{s}{s^2+3s-4} \right \}=\frac{1}{5}e^{t}+\frac{4}{5}e^{-4t}